# MIST: Mutual Information Estimation via Supervised Training

### v1

@@ 4 Experiments
Empirically, we find that our learned estimators outperform existing methods across sample sizes and dataset dimensions, particularly in low-sample settings (4.1). We provide an empirical study of the behavior of estimators under increasing dimensionality, finding that on average, our models require roughly half the number of samples as the best baseline to provide reliable MI estimates (4.2). The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives (4.3), while inference is orders of magnitude faster than existing neural methods (4.4). Finally, the learned estimators generalize to unseen distributions and sample sizes (4.5), and benefits from the mix of dimensions and size of the training dataset (4.6).
4.1 Comparison with Existing Estimators
We begin by benchmarking our estimators against existing methods on a diverse set of synthetic distributions. In Table 1, we present the average MSE loss of each estimator on $\mathcal{M}_{\text{test}}$, grouped by ranges of meta-datapoint size for clarity. We use the smaller test set of 2,340 meta-datapoints to evaluate all baselines.
@@ 4.1 Comparison with Existing Estimators
Table 1: Comparison of MI estimation baselines and our proposed methods across different distribution families (IMD and OoMD) and varying numbers of samples per meta-datapoint on the $\mathcal{M}_{\text{test}}$. Our estimators substantially outperform existing methods across sample size and distribution settings, including for distributions unseen during training. Shown are the averaged MSE loss values (mean ± stddev) of each estimator, grouped by the number of samples $n$ given to the estimator. The standard deviation is computed from 100 bootstrap resamples for each regime.
In this challenging regime of low sample sizes, varying dimensionality, and non-Gaussian distributions, recent estimators struggle to outperform the classical KSG baseline. This aligns with recent findings demonstrating that evaluation on more complex settings reveals weaknesses in modern estimators (bmi_benchmark; lee2024a).
Our learned estimators substantially outperform all baselines. Relative to KSG (the next strongest method), our models achieve approximately $10\times$ lower error on distributions seen during training (IMD), and $\sim 5\times$ lower error on unseen distributions (OoMD). A breakdown by distribution family is provided in Appendix C.3. For subsequent large-scale evaluation, we focus on KSG as the baseline, since it is both computationally feasible on the extended test set and remains the strongest baseline estimator in Table 1, particularly in low-sample regimes.
@@
Figure 2: Heatmaps providing a detailed analysis of the performance of the two proposed MI estimators and the KSG baseline. The data dimensions are grouped into three categories, and within each group, meta-datapoints are sorted by the number of samples. Shown are the average values of MSE, bias, variance, and confidence interval (CI cov.) width across $\mathcal{M}_{\text{test-extended}}$. CI coverage denotes the fraction of samples per group for which the true MI lies within the 95% bootstrap confidence interval.
@@
4.2 Scaling Behavior
Traditional evaluations of MI estimators have been constrained to relatively simple distributions and high-sample regimes. While recent efforts such as the BMI benchmark have enabled more comprehensive studies on complex distributions, systematic empirical analysis of scaling behavior has remained limited, largely due to the computational expense of evaluating existing methods. Leveraging our larger test set $\mathcal{M}_{\text{test-extended}}$, we conduct a detailed empirical study of estimator performance across a wide range of dimensions and sample sizes. This enables us to characterize performance frontiers in terms of data requirements. This analysis addresses an important question for practitioners: How many samples are needed for a reliable estimate?
@@ 4.2 Scaling Behavior
4.3 Calibration
@@
4.4 Inference Efficiency
@@
We estimated the inference time of each baseline on the entirety of $\mathcal{M}_{\text{test-extended}}$ and report the results in Appendix C.4. The high computational costs (>> 1 week) of evaluating all baselines except for CCA and KSG on $\mathcal{M}_{\text{test-extended}}$, along with the poor performance measured for CCA on the $\mathcal{M}_{\text{test}}$ for low sample sizes, further motivate our selection of KSG as the only baseline for other experiments in this work.
Figure 6: Inference time for each method as a function of the sample size. Results are averaged over 100 inferences (5 for MINDE), each performed over 10 meta-datapoints.
4.5 Generalization to Unseen Distributions, Dimensions, and Sample Sizes
The previous sections demonstrate that the learned algorithm generalizes effectively to unseen distribution families, which is evaluated at dimensions and sample sizes seen during training. We find that by learning to infer MI across diverse distributions and sample regimes, our estimators learn appropriate regularization strategies that transfer even to distribution families not encountered during training. The strength of these results motivates a bolder question: if meta-learned MI algorithms can generalize to unseen distributions, can they also generalize to these distributions at dimensions and sample sizes beyond their training experience?
We conduct a controlled study by training models on restricted subsets of the training data, specifically training on meta-datapoints in $M_{\text{train}}$ with a dimensionality $D<16$ and a sample size $S<300$. We evaluate performance on our full test meta-dataset $\mathcal{M}_{\text{test-extended}}$, which therefore includes distributions with higher dimensionalities and sample sizes and new distributions. Figure 7 presents two loss heatmaps, one for $\mathcal{M}_{\text{test-extended,IMD}}$ and one for $\mathcal{M}_{\text{test-extended,OoMD}}$.
We find that our estimators generalize well to new dimensions and sample sizes for distributions that are in the training meta-distribution (IMD heatmap). Similarly, we find that our models generalize well to new distributions when the data dimensionality is similar to that during training (OoMD heatmap, left half). However, they struggle to generalize to unseen distributions when the data also has higher dimensionality than any seen during training (OoMD heatmap, right half).
@@ 4.5 Generalization to Unseen Distributions, Dimensions, and Sample Sizes
4.6 Impact of Variable Input Dimensions
A natural question that arises from our approach is the impact of including multiple dimensions in model training. Of course, training on a mix of dimensions allows the development of a general-purpose estimator that can achieve high performance across settings, as shown in the above experiments. However, there may be settings where practitioners know their data dimension apriori, and may consider producing a training dataset of only that dimension. To study these questions, we trained separate (specialized) copies of MIST on subsets of $\mathcal{M}_{\text{train}}$ filtered to contain only meta-datapoints with a fixed input dimension (e.g., 4, 8, 16 or 32). This filtering procedure reduces the number of training data points substantially (by a factor of 31). To provide a comparison with the same number of samples, we also trained an additional (reduced data) MIST model on a uniformly down-sampled version of $\mathcal{M}_{\text{train}}$ to match the training size of the individual dimension models. By comparing the base MIST (full training set) with the specialized models (only one dimension), we can measure the impact of scaling both the dimensions and dataset size. By comparing the reduced data model (downsampled train set) and the specialized models (only one dimension), we can isolate the impact of training on multiple dimensions, without increasing dataset scale. In Figure 8, we provide these comparisons for each input dimension in $\mathcal{M}_{\text{test-extended}}$.
@@ 5 Discussion
Trade-offs and limitations. By embracing an empirical learning framework, we inherit the usual risks of statistical learning. The learned estimator’s accuracy depends on the diversity and coverage of the meta-training distributions, and its generalization to distant out-of-distribution families cannot be guaranteed a priori. Our estimator’s behavior is learned, rather than designed, and therefore does not carry the theoretical consistency guarantees of classical MI estimators—namely, that the estimate converges to the true MI as $n\to\infty$. We studied our estimator’s behavior on self-consistency tests from MINE_exp_variance_consistency_tests in Appendix D, and find that while our estimators are approximately self-consistent under the independence test, they are not self-consistent and produce high variance prediction ratios for the data processing and additivity tests. We believe that a learning-based MI framework offers opportunity for designing models, datasets, and losses to induce such desired behavior characteristics, but we believe it is important to recognize that our estimators do not carry these properties innately. In high-sample settings, our learned estimator cannot directly leverage all available samples during one inference pass and other estimators may be preferred if time and compute resources are available. Moreover, learned estimators are not inherently interpretable and may result in unforeseen failure modes. These limitations parallel the general challenges faced by machine learning models in other domains, highlighting the importance of careful validation and uncertainty quantification.
Controllability and adaptability. A key strength of this framework lies in its controllability. Our estimators’ behavior are largely determined by the synthetic datasets created; practitioners can customize, interrogate, and design datasets to improve estimator quality and reliability. Since mutual information is invariant under invertible transformations, our meta-dataset can be mapped to a target modality using normalizing flows without altering its ground-truth MI. This property enables the generation of diverse synthetic meta-datasets that reflect the characteristics of arbitrary data domains, such as images, text, tabular data, or multimodal signals. By tailoring the training data through such transformations, one can design estimators specialized for particular classes of distributions or downstream applications. This mechanism opens the door to what may be viewed as a foundational model for information-theoretic estimation, one that can be adapted via fine-tuning to specific tasks and modalities.
@@ A.1 Definitions of Mutual Information
A commonly used expression of MI is in terms of the Kullback-Leibler (KL) divergence between the joint distribution and the product of marginals (donsker_varadhan_kl_div_mi_definition):
@@ A.2 MI Estimators
We provide an overview of prominent MI estimators for reference in Table 2, categorized by their prediction category, the MI definition considered, and the estimation task. Previous works such as MINE_exp_variance_consistency_tests have described variational estimators as either discriminative or generative, with discriminative methods learning the ratio between joint and marginal densities directly and generative methods estimating the densities separately before computing the density ratio. We broaden this categorization into density and density ratio estimators.
Method Category $I(X;Y)$ Estimation Task
@@
KSG Density =H(X)+H(Y)H(X,Y)=H(X)+H(Y)-H(X,Y) Implicitly estimate H()H(*) by writing H()H(*) in terms of KNN distances
@@
MIENF Density $I(f_{X}(X);f_{Y}(Y))$ Learn flows $f_{X}$, $f_{Y}$ to Gaussian marginals
LMI Density $\approx I(Z_{x};Z_{y})$ Learn neural networks for embeddings $Z_{x}=g_{X}(X)$, $Z_{y}=g_{Y}(Y)$, then apply KSG
@@
MINDE Density $D_{KL}(p(x,y)\|p(x)p(y))$ Learn scores $\nabla\log p(x,y)$, $\nabla\log p(x)$ $\nabla\log p(y)$
@@
MINE Density ratio $\geq\sup_{\theta}\mathbb{E}_{P_{XY}}[T_{\theta}]-\log(\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}}])$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approx. $\log(\frac{p(x,y)}{p(x)p(y)})$
NWJ Density ratio $\geq\sup_{\theta}\mathbb{E}_{P_{XY}}[T_{\theta}]-\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}-1}]$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approx. $\log\frac{p(x,y)}{p(x)p(y)}$
InfoNCE Density ratio $\geq\mathbb{E}_{(x,y)\sim P_{n}}\left[\log\frac{f_{\theta}(x,y)}{\frac{1}{n}\sum_{j=1}^{n}f_{\theta}(x,y_{j})}\right]$ Learn classifier critic $f_{\theta}(x,y)\propto\frac{p(x,y)}{p(x)p(y)}$
SMILE Density ratio $\geq\mathbb{E}_{P}[T_{\theta}]-\log\mathbb{E}_{Q}[\text{clip}{\tau}(e^{T_{\theta}})]$ * Learn classifier critic $f_{\theta}(x,y)$ to maximize bound.
MIST From Samples $\approx F(x,y)$ Learn estimator $F$ by direct optimization
$\text{clip}_{\tau}(z)\in[e^{-\tau},e^{\tau}]$.

### v3

@@ 4 Experiments
Our experiments demonstrate that MIST outperforms existing methods across sample sizes and dimensions, particularly in low-sample regimes (4.1). MIST exhibits lower bias, lower variance, and higher CI coverage with increasing dimensions and samples (4.2). On average, MIST requires half the samples of the best baseline for reliable estimates in higher dimensions (4.3). The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives (4.4), while being orders of magnitude faster than existing neural methods (4.6). MIST generalizes to unseen distributions and sample sizes (4.5) and benefits from training on diverse dimensions and dataset sizes (4.7).
4.1 MIST Outperforms Existing Estimators, Even on Unseen Distributions
We begin by benchmarking our estimators against existing methods on a diverse set of synthetic distributions. In Table 1, we present the average MSE loss of each estimator on $\mathcal{M}_{\text{test}}$, grouped by ranges of meta-datapoint size for clarity. We use the smaller test set of 2,340 meta-datapoints to evaluate all baselines. In this challenging regime of low sample sizes, varying dimensionality, and non-Gaussian distributions, recent estimators struggle to outperform the classical KSG baseline. This aligns with recent findings demonstrating that evaluation on more complex settings reveals weaknesses in modern estimators (Czyż et al., 2023; Lee and Rhee, 2024). Our learned estimators substantially outperform all baselines. Relative to KSG (the next strongest method), our models achieve approximately $10\times$ lower error on distributions seen during training (IMD), and $\sim 5\times$ lower error on unseen distributions (OoMD). A breakdown by distribution family is provided in Appendix C.3. For subsequent large-scale evaluation, we focus on KSG as the baseline, since it is both computationally feasible on the extended test set and remains the strongest baseline estimator in Table 1, particularly in low-sample regimes.
Table 1: MIST estimators substantially outperform existing methods across sample size ($n$) and distributions, including for distributions unseen during training (OoMD, right). Shown are the averaged MSE loss values (mean ± stddev) of each estimator, grouped by the number of samples $n$ given to the estimator, and using 100 bootstrap samples for the standard deviation.
@@ 4.1 MIST Outperforms Existing Estimators, Even on Unseen Distributions
InfoNet $3.2\mathrm{e}^{1}$ $\pm 9.3\mathrm{e}^{1}$ $3.5\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$ $3.7\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$ $3.9\mathrm{e}^{1}$ $\pm 8.2\mathrm{e}^{1}$ $4.2\mathrm{e}^{1}$ $\pm 9.9\mathrm{e}^{1}$ $5.5\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$
FMMI $2.8\mathrm{e}^{1}$ $\pm 8.4\mathrm{e}^{1}$ $2.5\mathrm{e}^{1}$ $\pm 9.8\mathrm{e}^{1}$ $1.7\mathrm{e}^{1}$ $\pm 6.5\mathrm{e}^{1}$ $3.3\mathrm{e}^{1}$ $\pm 7.4\mathrm{e}^{1}$ $2.8\mathrm{e}^{1}$ $\pm 7.7\mathrm{e}^{1}$ $3.1\mathrm{e}^{1}$ $\pm 8.2\mathrm{e}^{1}$
VCE $1.5\mathrm{e}^{1}$ $\pm 3.9\mathrm{e}^{1}*$ $1.4\mathrm{e}^{1}$ $\pm 4.3\mathrm{e}^{1}$ $1.6\mathrm{e}^{1}$ $\pm 3.6\mathrm{e}^{1}$ $2.6\mathrm{e}^{1}$ $\pm 6.5\mathrm{e}^{1}*$ $2.4\mathrm{e}^{1}$ $\pm 6.6\mathrm{e}^{1}$ $2.4\mathrm{e}^{1}$ $\pm 5.9\mathrm{e}^{1}$
@@
* VCE failed to produce an estimate for 99/216 points (118/252) in the IMD (OoMD) subsets.
4.2 MIST Exhibits Lower Bias and Variance Across Settings
@@
Figure 3: MIST estimators widely outperform KSG, remaining substantially less biased and lower variance across dimensions groups and sample sizes in $\mathcal{M}_{\text{test-extended}}$. CI coverage denotes the fraction of samples per group for which the true MI lies within the 95% bootstrap confidence interval.
4.3 MIST Requires Fewer Samples to Scale to Higher Dimensions
Traditional evaluations of MI estimators have been constrained to relatively simple distributions and high-sample regimes. While recent efforts have enabled more comprehensive studies on complex distributions, systematic empirical analysis of scaling behavior has remained limited, largely due to the computational expense of evaluating existing methods. Leveraging our larger test set $\mathcal{M}_{\text{test-extended}}$, we conduct a detailed empirical study of estimator performance across a wide range of dimensions and sample sizes. This enables us to characterize performance frontiers in terms of data requirements. This analysis addresses an important question for practitioners: How many samples are needed for a reliable estimate?
@@ 4.3 MIST Requires Fewer Samples to Scale to Higher Dimensions
4.4 MIST Quantile-Based Intervals are More Reliable Than Bootstrapping
@@
4.5 MIST Generalizes to Unseen Distributions, Dimensions, and Sample Sizes
The previous sections demonstrate the unique ability of meta-learned estimators to generalize to unseen distribution families (OoMD) at test-time. The strength of these results motivates a bolder question: if meta-learned MI algorithms can generalize to unseen distributions, can they also generalize to dimensions and sample sizes beyond their training experience?
We conduct a controlled study by training models on only the subset of $M_{\text{train}}$ with $D<16$ and $n<300$, then evaluating on the full test set $\mathcal{M}_{\text{test-extended}}$ which includes unseen higher dimensions ($D>16)$, unseen larger sample sizes ($n>300$), and unseen new distributions (OoMD). Figure 6 presents loss heatmaps for in-distribution (IMD) and out-of-distribution (OoMD) test sets. We find that our estimators generalize well to new dimensions and sample sizes for distributions that are in the training meta-distribution (IMD heatmap). Similarly, we find that MIST models generalize well to new distributions when the data dimensionality is similar to that during training (OoMD heatmap, left half). However, they struggle to generalize to unseen distributions when the data also has higher dimensionality than any seen during training (OoMD heatmap, right half).
@@ 4.5 MIST Generalizes to Unseen Distributions, Dimensions, and Sample Sizes
4.6 MIST Achieves Large Gains in Inference Efficiency
Figure 7: Average inference time for each method with increasing sample size.
@@ 4.6 MIST Achieves Large Gains in Inference Efficiency
The combination of fast inference and efficient uncertainty quantification (as discussed in Section 4.4) makes our approach practical for large-scale applications where repeated MI estimation is required, such as hyperparameter tuning, feature selection, or iterative model training (see Appendix E for an example of use in feature selection). We estimated the inference time of each baseline on the entirety of $\mathcal{M}_{\text{test-extended}}$ and report the results in Appendix C.4. The high computational costs (>> 1 week) of evaluating all baselines except for CCA and KSG on $\mathcal{M}_{\text{test-extended}}$, along with the poor performance measured for CCA on the $\mathcal{M}_{\text{test}}$ for low sample sizes, further motivate our selection of KSG as the only baseline for other experiments in this work. On our multidimensional test set, InfoNet’s slicing overhead offsets its efficiency gains over optimized methods like MINE.
4.7 Training on Multiple Dimensions Benefits High-Dimensional Estimation
In contrast to existing paradigms, MIST models learn over a diverse meta-distribution rather than fitting an individual dataset of interest. This approach opens many relevant empirical questions regarding the design of the meta-distribution. One of the most prominent design choices is the choice to train over many dimensions or to train a distribution-specific model. While training over many dimensions has the benefit of generalized use, practitioners may know the dimension of their data a priori.
@@ 4.7 Training on Multiple Dimensions Benefits High-Dimensional Estimation
To study these questions, we trained separate (specialized) MIST models on subsets of $\mathcal{M}_{\text{train}}$ filtered to contain only meta-datapoints with a fixed input dimension (e.g., 4, 8, 16 or 32). This filtering procedure reduces the number of training data points substantially (by a factor of 31). To provide a comparison with the same number of samples, we also trained additional (reduced data) MIST models on a uniformly down-sampled version of $\mathcal{M}_{\text{train}}$ to match the number of samples seen by the individual dimension models. By comparing the base MIST (full training set) with the specialized models (only one dimension), we can measure the impact of scaling both the dimensions and dataset size. By comparing the reduced data model (downsampled train set) and the specialized models (only one dimension), we can isolate the impact of training on multiple dimensions, without increasing dataset scale. In Figure 8, we provide these comparisons for each input dimension in $\mathcal{M}_{\text{test-extended}}$.
@@ 5 Discussion
Trade-offs and limitations. By embracing an empirical learning framework, we inherit the usual risks of statistical learning. The learned estimator’s accuracy depends on the diversity and coverage of the meta-training distributions, and its generalization to distant out-of-distribution families cannot be guaranteed a priori. Our estimator’s behavior is learned, rather than designed, and therefore does not carry the theoretical consistency guarantees of classical MI estimators—namely, that the estimate converges to the true MI as $n\to\infty$. We studied our estimator’s behavior on self-consistency tests from Song and Ermon (2019) in Appendix D, and find that while our estimators are approximately self-consistent under the independence test, they are not self-consistent and produce high variance prediction ratios for the data processing and additivity tests. In Appendix E, we study the utility of MIST predictions downstream in a task of learned feature selection, using MINE as a baseline. In this setting, we find that MIST-guided learning was more accurate, lower-variance, and more stable across a range of sample sizes and noise levels. A learning-based MI framework offers opportunity for designing models, datasets, and losses to induce desired behavior characteristics, but we believe it is important to recognize that our estimators do not carry these properties innately and empirical evaluation is required. Another important consequence of MIST being a learned estimator is that the range of MI values predicted by MIST are limited to those seen during training. In high-sample settings, our learned estimator cannot directly leverage all available samples during one inference pass and other estimators may be preferred if time and compute resources are available. Moreover, learned estimators are not inherently interpretable and may result in unforeseen failure modes. These limitations parallel the general challenges faced by machine learning models in other domains, highlighting the importance of careful validation and uncertainty quantification.
[…]
@@ A.1 Definitions of Mutual Information
A commonly used expression of MI is in terms of the Kullback-Leibler (KL) divergence between the joint distribution and the product of marginals (Donsker and Varadhan, 1975):
@@ A.2 MI Estimators
We provide an overview of prominent MI estimators for reference in Table 2, categorized by the type of estimation they perform, the MI definition considered, and the specific estimation task. Previous works have often adopted the terminology introduced by Song and Ermon (2019). This framework categorizes variational estimators as either discriminative or generative: generative methods estimate the densities individually and discriminative methods estimate the ratio between joint and marginal densities.
Given the diversity of recent estimation approaches, we broaden this categorization into density and density ratio estimators. In this categorization, density estimators estimate MI by approximating the form or properties of densities $P_{X},P_{Y},\text{and }P_{X,Y}$. This includes learning the densities explicitly with a generative model (GM, KDE, VCE), approximating densities through k-nearest-neighbor distances (KSG), learning flows to transform the marginal distributions to a tractable distribution (MIENF), or learning score functions (MINDE). In contrast, density ratio estimators estimate MI by modeling the relationship between densities, primarily in the form of the density ratio $log(\frac{P_{X,Y}}{P_{X}P_{Y}})$, without modeling the densities individually. This estimation can be performed by learning classifiers (’critics’) to distinguish joint and marginal samples (MINE, SMILE, InfoNCE), or learning coupling transforms that map between the joint and marginal product (FMMI).
We chose this naming convention to better capture the diversity of existing methods and highlight the core task performed by each method. We hope that this naming convention can help avoid reader confusion between the modeling framework and the core MI estimation task. For example, works like FMMI are generative in terms of modeling choices (flows), but discriminative in terms of the MI estimation task. Similarly, many classical estimators are not generative in terms of modeling choices (binning, KSG), but are generative in terms of MI estimation. To improve clarity, we group methods by their estimation task (estimate a density, or a density ratio), regardless of the approach used to perform the estimation. Still, this categorization, as with any abstraction, removes some nuance. We provide this appendix section as a resource for those looking to gain a better holistic understanding of existing approaches, and encourage readers to read the respective papers for precise details.
Method Type $I(X;Y)$ Estimation Task
@@
KSG Density =H(X)+H(Y)H(X,Y)=H(X)+H(Y)-H(X,Y) Implicitly estimate $H(\cdot)$ by writing MI’s entropy definition in terms of KNN distances
@@
MIENF Density $\approx I(f_{X}(X);f_{Y}(Y))$ Learn flows $f_{X}$, $f_{Y}$ to Gaussian marginals
LMI Density $\approx I_{KSG}(Z_{x};Z_{y})$ Learn embeddings $Z_{x}=g_{X}(X)$, $Z_{y}=g_{Y}(Y)$ via cross-predictive autoencoders $g_{X},g_{Y}$, then apply KSG
@@
MINDE Density $\begin{aligned} &=D_{KL}(p(x,y)\|p(x)p(y))\\ &\approx\|\nabla\log p(x,y)-\nabla\log p(x)p(y)\|^{2}\end{aligned}$ Learn scores $\nabla\log p(x,y)$, $\nabla\log p(x)$, $\nabla\log p(y)$ with a diffusion model
@@
MINE Density Ratio $\geq\sup_{\theta}\mathbb{E}_{P_{X,Y}}[T_{\theta}]-\log(\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}}])$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approximating $\log(\frac{p(x,y)}{p(x)p(y)})$
NWJ Density Ratio $\geq\sup_{\theta}\mathbb{E}_{P_{X,Y}}[T_{\theta}]-\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}-1}]$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approximating $\log(\frac{p(x,y)}{p(x)p(y)})+1$
InfoNCE Density Ratio $\geq\mathbb{E}_{(x,y)\sim P_{X,Y}}\left[\log\frac{f_{\theta}(x,y)}{\frac{1}{n}\sum_{j=1}^{n}f_{\theta}(x,y_{j})}\right]$ Learn classifier critic $f_{\theta}(x,y)\propto\frac{p(x,y)}{p(x)p(y)}$
SMILE Density Ratio $\geq\mathbb{E}_{P_{X,Y}}[T_{\theta}]-\log\mathbb{E}_{P_{X}\otimes P_{Y}}[\text{clip}_{\tau}(e^{T_{\theta}})]$ * Learn classifier critic $T_{\theta}(x,y)$ to maximize bound, approximating $\log(\frac{p(x,y)}{p(x)p(y)})$
FMMI** Density Ratio $\begin{aligned} &=-\mathbb{E}_{T\sim\mathcal{U}[0,1]}[\text{div }v(Z_{T},T)]\\ &Z_{0}\sim P_{X}\otimes P_{Y},\,Z_{1}\sim P_{X,Y}\end{aligned}$ Train velocity network $v(z,t)$ via flow matching, estimate MI as negative expected divergence of $v$
InfoNet Density Ratio $\geq\sup_{\theta}\mathbb{E}_{P_{X,Y}}[\theta]-\log(\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{\theta}])$ Pretrain network $F(x,y)\rightarrow\theta$, where $\theta$ is MINE’s optimal discriminant as as a look-up table
MIST Meta- Learned $\approx F(x,y)$ Learn estimator $F$ by direct optimization
* $\text{clip}_{\tau}(z)\in[e^{-\tau},e^{\tau}]$.
** FMMI learns a coupling transform between joint and marginal distributions. This estimation implicitly encodes the density ratio, which could be derived from the velocity field with an integration. However, FMMI does not perform this integral and does not estimate the density ratio at each point. Instead, FMMI estimates MI as an expectation of the field’s divergence. FMMI authors describe their method as a hybrid method. We categorize FMMI as a density ratio method due to the focus on learning the relationship between joint and marginal distributions, rather than modeling the densities individually.